A thin large size active diamond target for a dark photon search experiment

R. Assiro^{1} ,A. Caricato^{1}^{,}^{2} ,G. Chiodini^{1} ,M. Corrado^{2} ,M.De Feudis^{1}^{,}^{2} ,G. Fiore^{1} ,M. Martino^{1}^{,}^{2} ,A. Miccoli^{1} ,F. Oliva^{1}^{,}^{2} ,G. Maruccio^{2}^{,}^{3} ,A.G. Monteduro^{3} ^{,}^{2} ,C. Pinto^{1}^{,}^{2} and S. Spagnolo^{1}^{,}^{2}.

1Istituto Nazionale di Fisica Nucleare, Sezione di Lecce, Italy.

2Dip. di Matematica e Fisica “Ennio de Giorgi”, Università del Salento, Italy.

3CNR NANOTEC di Lecce, Italy.

The Lecce group is responsible of the design, construction and operation of the diamond active target of PADME, a new experiment proposed to search for a dark photon in the annihilation of e^{+}e^{−} at INFN Laboratori Nazionali di Frascati.

The compelling cosmological evidence for dark matter is still missing an experimental observation at microscopic level of non-ordinary particles representing elementary constituents of the dark matter. This might be the natural consequence of scenarios where the new particles do not interact directly with ordinary matter but manifest themselves only through portal fields connecting our world with new hidden sectors.

A theoretically simple and well motivated model assumes the existence of an abelian gauge symmetry U(1), in addition to the Standard Model gauge symmetry group, whose gauge boson, the dark photon U, couples directly to the matter fields of the hidden sectors and mixes to the Standard Model (SM) photon thus providing a connection with the visible sector.

Dark photons were searched in accelerator experiments by looking for resonant fermion pair production in meson decays from beam dumps, or in e+e- collisions, but no evidences were found so far [1].

At Laboratori Nazionali di Frascati an alternative approach is proposed for the new experiment PADME. At the Beam Test Facility (BTF), the invisible decay of a dark photon will be searched in positron-electron annihilations by triggering on single photon final states and reconstructing the missing mass of the event [2].

The PADME detector should be built in 2017. The accessible dark photon mass range is from 1 MeV to 23.7 MeV, for the present Frascati linac maximum positron energy of 550 MeV. A data taking campaign of about six months in 2018 should ensure exclusion limits for the dark photon coupling ϵ of about 0.001, for a bunch length of 40 ns (corresponding to 10^{13} e^{+} on target), or about 0.0005, for a bunch length of 480 ns (corresponding to 10^{14} e^{+} on target).

The sensitivity of the experiment is enought to confirm or reject the hypothesis that the discrepancy between the expected and measured value of the muon anomalous magnetic moment (more than a 3 sigma effect) is due to the dark photon. In fact, the anomaly could be explains by a new vector interaction in addition to the photon interaction. The standard model g-2 correction due to photon exchange is about \frac{α_{e}}{4π} ≈10^{−3} for both electron and muon. A dark photon with a mass much smaller than the muon mass and much higher than electron mass, should correct the muon g-2, but not the electron g-2. In addition, the g-2 correction due to the dark photon mixing should be about Δ(g − 2)_{μ} ≈10^{−9} ≈ \frac{α_{e}}{4π} ϵ^{2} ≈10^{−3}ϵ^{2}, which means ϵ ≈10^{−3}, this is in the PADME physics reach.

In order to improve the resolution in reconstructing the missing mass, the experiment PADME measures bunch-by-bunch the beam position and profile by the carbon nuclear target. The adopted solution is to build a thin diamond active target having a spatial resolution better than 1 mm in the two directions X and Y transverse to the beam.

Our group in Lecce proposed to build as active target of PADME a polycrystalline diamond sensor fabricating electrods not by the traditional metal multi-layers, but by the method developed in Lecce, which makes graphitic contacts on diamonds by excimer laser [3]. The advantages are two folds: first, the target is full carbon without metals contamination and, second, the damage probability due to manipolation is reduced. In fact, the target thickness should be between 50 and 100 μm, to minimize the photon re-interaction, and the size about 2x2cm^{2}, to fully intercept the positron beam. Diamond sensors such thin and large were never used so far and they are expected to be quite fragile to work with.

The active target is segmented on both sides with 1 mm orthogonal strips. The choice of strip geometry and not pixel geometry is related to Figure 1. First full carbon PADME active target prototype realized in Lecce.

the unacceptable material introduced in the active area for the last solution. A consequence of strip geometry is that the active target can reconstruct only X and Y profile independently and not a two dimensional X-Y profile, which is possible only with pixel geometry or with an out-of-beam image device (for example Cherenkov image device). This limitation should not be a problem for our application because the beam shape should be known.

With respect to diamond detector “state of the art” the most challenging aspect of an active diamond target for the PADME experiment is related to the small thickness and large area, as it was said before, their make the sensor quite fragile. Instead, the quite low signal (CCD < thickness/2) for a MIP particle is likely not a problem for PADME, because the bunch of electrons is composed of few thousand particles. Nevertheless, response uniformity, gain fluctuations and linearity must be verified.

In October 2015, a full size active target diamond propotype was build in Lecce with XY strips on both faces having a pitch of 1 mm by an UV ArF excimer laser (see Fig. 1). The prototype was tested in November 2015 at the Beam Test Facility (BTF) at INFN Laboratori Nazionali di Frascati using e^{+}e^{−} beams of about 450 MeV energy. The preliminary results show a very good spatial resolution of about 0.2 mm on the average X and Y beam position for bunch multiplicity of about 10,000 particles (see Fig. 1).

REFERENCES

1. Raggi M. and Kozhuharov V. “Results and perspectives in dark photon physics”, Rivista del nuovo cimento. Vol. 38, N. 10, 449-505 Figure 2. Distribution of the BTF beam X position estimated by the charge weighting method applied to the PADME diamond active target prototype realized in Lecce.

(2015).

2. Raggi M. and Kozhuharov V. “Proposal to Search for a Dark Photon in Positron on Target Collisions at DAΦNE Linac”, Advances in High Energy Physics ID 959802 (2014).

3. M. De Feudis et al.“Realization and characterization of graphitic contacts on diamond by means of laser”, “IV National Workshop on Plasmi Sorgenti Biofisica e Applicazioni (PSBA)”, Lecce (Italy), 17-18 Oct.2014.